In a right triangle, if one acute angle has measure 6x-10 and the other has measure 2x-4, find the larger of these two angles.
A) 13 degrees
B) 22 degrees
C) 47 degrees
D) 68 degrees

Answer by stanbon(48535)

In a right triangle, if one acute angle has measure 6x-10 and the other has measure 2x-4, find the larger of these two angles.
A) 13 degrees
B) 22 degrees
C) 47 degrees
D) 68 degrees
6x-10+2x-4=90
8x-14=90
8x=104
x=13
Larger angle is 6x-10=6(13)-10=68 degrees
Cheers,
Stan H.